OpenStudy (anonymous):

A constant term is: Select one: a. A monomial with a variable. b. A monomial without a variable. c. A binomial with a variable. d. A binomial without a variable.

OpenStudy (theedwardsfamily):

this will help u: http://en.wikipedia.org/wiki/Constant_term

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

or D

OpenStudy (theedwardsfamily):

A

OpenStudy (anonymous):

it doesnt have a variable tho

OpenStudy (anonymous):

so how is it A?

OpenStudy (anonymous):

No Is A

OpenStudy (anonymous):

oh .. sorry my mistake then

OpenStudy (anonymous):

i think D

OpenStudy (theedwardsfamily):

did you look at the link? @denisse18

OpenStudy (anonymous):

yes but do not have the information necessary

OpenStudy (anonymous):

well according to this term has a constant non-variable

OpenStudy (anonymous):

see thats what i said

OpenStudy (anonymous):

i think its D too

OpenStudy (anonymous):

because they said it was polynomial so i guess bi passes as poly

OpenStudy (anonymous):

well idk its your choice A ( like edward said) or D

OpenStudy (anonymous):

A or D ?

OpenStudy (anonymous):

id go with D ... because what i looked up says its polynomial and that it doesnt have variables

OpenStudy (anonymous):

"The term in a simplified algebraic expression or equation which contains no variable(s). If there is no such term, the constant term is 0."

OpenStudy (anonymous):

"The Constant Term in an expression or equation has a fixed value and does not contain variables."

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

see i was right :D

OpenStudy (anonymous):

thanks for the medal merp :)

OpenStudy (anonymous):

Can you help me another question??

OpenStudy (anonymous):

i can try ..

OpenStudy (anonymous):

but hurry i only have 10 min

OpenStudy (anonymous):

Write the following function F as a composition of 2 functions: F(x) = |3x2 + 2x – 8| Select one: a. f(x) = |x| and g(x) = 3x2 + 2x – 8 b. g(x) = |x| and f(x) = 3x2 + 2x – 8 c. f(x) = x and g(x) = |3x2 + 2x – 8| d. Both A and B

OpenStudy (anonymous):

i cant see it :/ llike a whole bunch of question marks are showing up

OpenStudy (anonymous):

but with functions you have to pug it in

OpenStudy (anonymous):

��������������������������� <---- these are showing up

OpenStudy (anonymous):

mmm Write the following function F as a composition of 2 functions: F(x) = |3x2 + 2x – 8| Select one: a. f(x) = |x| and g(x) = 3x2 + 2x – 8 b. g(x) = |x| and f(x) = 3x2 + 2x – 8 c. f(x) = x and g(x) = |3x2 + 2x – 8| d. Both A and B

OpenStudy (anonymous):

i still cant see it :/

OpenStudy (anonymous):

but i have to go sorry i couldnt help idk whats wrong with openstudy .. Good Luck !